OpenStudy (anonymous):

mike wants to make meatloaf. his recipe uses a total of 7 pounds of meat. if he uses 3 to 1 ratio of beef to pork how much pork will he use? enter your answer as a mixed number in simplest terms

OpenStudy (anonymous):

?

OpenStudy (shinalcantara):

let 'x' be the amount of pork 'y' be the amount of beef x + y = 7 ----equation 1 knowing the ratio as 3:1 ( beef:pork) we can say that if in every 3x of beef there's x amount of pork then we'll have y=3x ---equation 2 substitute equation 2 to equation 1 x + y = 7 x + 3x = 7 4x = 7 x = 7/4 pounds of pork